Calculate the value of Kp for the equation

C(s)+CO2(g)−⇀2CO(g) Kp=?

given that at a certain temperature

C(s)+2H2O(g)−⇀CO2(g)+2H2(g) Kp1=3.69

H2(g)+CO2(g)−⇀H2O(g)+CO(g) Kp2=0.671
1
Expert's answer
2020-03-30T10:54:07-0400

Kp = Kp1 - 1 / Kp2 = 3.69 - 1 / 0.671 = 2.20